OpenStudy (anonymous):

NEED HELP WITH STANDARD FORM Now suppose you knew the volume of this object and the relation of the height to the radius, but did not know the radius. Rewriting the equation with one variable would result in a polynomial equation that you could solve to find the radius. Rewrite the formula using the variable x for the radius. Substitute the value of the volume found in step 2 for V and express the height of the of the object in terms of x plus or minus a constant. For example, if the height measurement is 4 inches longer than the radius, then the expression for the height will be (x + 4)

OpenStudy (anonymous):

V=Pi(r^2)(h), d=4, r=2, h=6, v= 75.36

OpenStudy (anonymous):

looks liek the instructions are written for you :P

OpenStudy (anonymous):

I don't understand the rewriting part

OpenStudy (anonymous):

Rewrite the formula using the variable x for the radius. V=Pi(r^2)(h) V=Pi((x)^2)(h) Substitute the value of the volume found in step 2 for V 75.36=Pi((x)^2)(h) i assume you mean v=75.36 from step 2? and express the height of the of the object in terms of x plus or minus a constant. 75.36=Pi((x)^2)(x+c) h=x+ c <- x plus minus a constant For example, if the height measurement is 4 inches longer than the radius, then the expression for the height will be (x + 4) 75.36=Pi((x)^2)(x+c)

OpenStudy (anonymous):

with me so far?

OpenStudy (anonymous):

okay yes, I got there !

OpenStudy (anonymous):

but how would I simplify that in standard form?

OpenStudy (anonymous):

you'd solve for x

OpenStudy (anonymous):

how? is this right: 314x^3+1256x^2-7536=0

OpenStudy (anonymous):

looks like you expanded right :) but i wanna try plugin in your givens.

OpenStudy (anonymous):

d=4, r=2, h=6, v= 75.36, h=x+ c 6=c+x sfc 6-x=c V=Pi(r^2)(h) 75.36=Pi((x)^2)(x+c) <- eq 1 so plugin in for c in eq1 75.36=Pi((x)^2)(x+(6-x)) 75.36-6 pi x^2 = 0 <- expanding and moving all terms to left hand side

OpenStudy (anonymous):

75.36-6 pi x^2 = 0 rearrange -6 pi x^2 + 75.36 = 0

OpenStudy (anonymous):

It all seems odd to me why you are given d=4, r=2, h=6, v= 75.36,

OpenStudy (anonymous):

ohh, which one would I use to Find the solutions to this equation algebraically using the Fundamental Theorem of Algebra, the Rational Root Theorem, Descartes' Rule of Signs, and the Factor Theorem?

OpenStudy (anonymous):

FTA i think.

OpenStudy (anonymous):

what's FTA?

OpenStudy (anonymous):

Fundamental Theorem of Algebra

OpenStudy (anonymous):

factor theorem looks right too !!

OpenStudy (anonymous):

okie dokie my polynomial is all I would need? Then I think I'm good from here. Thank you !

OpenStudy (anonymous):

still a little fishy but if you are cool with it :)